OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(2c) = square root of 5^(4c + 16) ? Select one: A. c = -5 over 2 B. c = 7 over 2 C. c = 4 D. c = -7 over 2

OpenStudy (jdoe0001):

check your equation, are the terms as \( 25^{2c}= \sqrt{5^{(4c+16)}} \) ? cuz I get a funky answer from that

OpenStudy (anonymous):

That's what the test says, this one got me too confused.

OpenStudy (jdoe0001):

ok, lemme check some more

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

I just found my error, ok, give a few secs

OpenStudy (anonymous):

Okay

OpenStudy (jdoe0001):

the one thing you have to bear in mind is that, whenever you want to get an EXPONENT DOWN, you use a Logarithm Cancellation rule, where the Log base is the same as the coefficient, lemme type it

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

$$ 25^{2c}= \sqrt{5^{(4c+16)}} \implies (5^2)^{2c}=\sqrt{5^{(4c+16)}}\\ \text{square both sides to get the expression OFF the root}\\ \pmatrix{(5^2)^{2c}}^2=\pmatrix{\sqrt{5^{(4c+16)}}}^2\\ 5^{8c}=5^{(4c+16)} $$

OpenStudy (jdoe0001):

$$ log_5{5^{8c}}=log_5{5^{(4c+16)}}\\ \text{log cancellation rule, base 5, to get the exponents down}\\ 8c=4c+16 $$ in this case both sides are of base 5, that is the NUMBER with the exponent is 5, thus, a log(5) gets both exponents

OpenStudy (jdoe0001):

and from there, I'd guess you know what to do :)

OpenStudy (rajee_sam):

Ans choice is C

OpenStudy (anonymous):

so then \[4c=16\] and then simplified it's \[c=4\] right?

OpenStudy (jdoe0001):

:D

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (jdoe0001):

yw